Ocean City Kite Company manufactures & sells kites for $6.00 each. The variable cost per kite is $3.50 with the current annual sales volume of 60,000 kites. This volume is currently Ocean City Kite's breaking even point. Use this information to determine the dollar amount of Ocean City Kite Company's fixed costs. (Round dollar value to the nearest whole dollar & enter as whole dollars only.)
Answer #1

Contribution margin=Sales-Variable cost  

=(6-3.5)=$2.5 per unit

Breakeven=Fixed expenses/Contribution margin

Hence total fixed expenses =(2.5*60,000)

=$150000